OpenStudy (anonymous):

A tower in Maintown sends radio signal a certain distance (in miles) according to the equation: x2 + y2 = 8,100. The Smith family lives 50 miles south and 70 miles east of Maintown. Part 1: Can the Smith family receive the radio signal from Maintown? (3 points)

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 50^2 + 70^2 = c^2 2500+4900 = c^2 keep going to solve for c this will give you the distance from Maintown to the Smith family home.

OpenStudy (anonymous):

Thank you very very much!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

I think the equation x2 + y2 = 8,100 is referring to the range. When you plug 50 and 70 in for x and y and the number is less than 8100 then they are in range, otherwise out of range. I don't interpret the equation the same way as @jim_thompson5910

OpenStudy (anonymous):

Okay i'll work it on! Thank you both! :)